A bone is found at an archaeological site and is placed into a mass spectrometer. It is found that there are 350 14C atoms for every 1
Answer #1

ratio of 14 C to 12C is

350 / 1*10^ 15 = 35*10^ -13

normal ration of 14 C to 12C is 1.2*10^-12.

The age of rock is

1.2*10^-12 / 35*10^-13 = e ^ -t/ T_1/2

0.34 = e ^ - t / 5730 y

t = 5730 Y